# The $B^{-}\to J/\psi\eta^{\prime}K^{-}$ reaction and the $Y(4710)$ and $K_{0}^{*}(1430)$ contributions

### v1

@@ Abstract
Motivated by the recent LHCb Collaboration analysis, we have investigated the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ by considering the contributions from the resonances $Y(4710)$ and $K_{0}^{*}(1430)$, where the $K_{0}^{*}(1430)$ is regarded as a dynamically generated state from the vector meson-vector meson interaction. Our results are in good agreement with the measured invariant mass distributions and the Dalitz plot of LHCb, which supports the molecular explanation of the $K_{0}^{*}(1430)$, and that the $Y(4710)$ can be produced in the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$. Since the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$ is Okubo-Zweig-Iizuka (OZI) suppressed for the charmonium state, the measurements of the $\mathcal{B}(Y(4710)\to J/\psi\eta^{\prime})$ through the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ or other processes is crucial to understanding the nature of the $Y(4710)$. Thus, we advocate that Belle II and LHCb Collaborations could perform the more precise analysis to confirm the evidence of the $Y(4710)$ state, which could be helpful to reduce the experimental uncertainties of its mass and width, and to explore its nature.
@@ I Introduction
Since the $X(3872)$ has been observed by the Belle Collaboration in 2003 [8], many charmonium-like states have been reported experimentally, and called many attentions [9, 10, 11, 12]. In 2023, the BESIII Collaboration observed a new enhancement structure $Y(4710)$ with a mass of $M=(4704.0\pm 52.3\pm 69.5)$ MeV and a width of $\Gamma=(183.2\pm 114.0\pm 96.1)$ MeV in the cross sections for the process $e^{+}e^{-}\to K_{S}^{0}K_{S}^{0}J/\psi$, and the statistical significance is $4.2\sigma$ [13]. Subsequently, using the data samples with an integrated luminosity of 5.85 $fb^{-1}$ collected at center-of-mass energies from 4.61 to 4.95 GeV, BESIII has measured the cross sections for the process $e^{+}e^{-}\to K^{+}K^{-}J/\psi$, and confirmed the resonance $Y(4710)$ with a significance over $5\sigma$ [14]. Its mass and width are measured to be $M=(4708^{+17}_{-15}\pm 21)$ MeV, $\Gamma=(126^{+27}_{-23}\pm 30)$ MeV, respectively [14], which are in consistent with the previous measurements of BESIII within the experimental uncertainties [13].
@@
Although the mass of $Y(4710)$ observed by BESIII is close to the $\psi(5S)$ mass (4704 MeV or 4711 MeV) predicted by the linear potential quark model [15, 16, 17], there are also different theoretical explanations for the structure of the $Y(4710)$ state [18, 19, 20]. In Refs. [21, 22], the authors deem that the $Y(4710)$ is a good candidate of the hidden-charm-hidden-strange tetraquark state $[sc]_{\tilde{V}}[\bar{sc}]_{A}-[sc]_{A}[\bar{sc}]_{\tilde{V}}$ with $J^{PC}=1^{--}$ within the framework of the QCD sum rules. In Ref. [23], the authors have studied the charmonium spectrum within an unquenched quark model including coupled channel effects, and found that the $Y(4710)$ can be interpreted as $\psi(4D)$ state. In addition, the $Y(4710)$ can also be interpreted as the hybrid candidates of $H_{1}(4507)$ [24] or $H_{1}[1^{--}](4812)$ [18]. It is notable that the mass and width of the $Y(4710)$ have very large uncertainties, and this state has not yet been listed in Review of Particle Physics (RPP) [25], which implies that the $Y(4710)$ is not well-understood, and still needs confirmation in more processes.
@@
The hadronic decays of the bottom hadrons provide an ideal platform to investigate the charmonium-like states [26, 35, 36, 29, 30, 31, 27, 32, 28, 34, 33]. Recently, the LHCb Collaboration has observed the process $B^{+}\to J/\psi\eta^{\prime}K^{+}$ [37], and given $\mathcal{B}(B^{+}\to J/\psi\eta^{\prime}K^{+})=(3.06\pm 0.29\pm 0.18\pm 0.04)\times 10^{-5}$. From Fig. 2(b) of Ref. [37], one can find a peak structure around 4.7 GeV in the $J/\psi\eta^{\prime}$ invariant mass distribution, close to the mass of $Y(4710)$. If this peak structure can be associated with $Y(4710)$, the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$, which is Okubo-Zweig-Iizuka (OZI) suppressed for the charmonium state decay [38, 39, 40], could be helpful to explore its nature, and the precise measurements of this process could be useful to determine the accurate properties of the $Y(4710)$.
@@
In the present work, we will investigate the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ [^1] In this work, we study the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$, the charge conjugation process of the $B^{+}\to J/\psi\eta^{\prime}K^{+}$, because it is convenient to deal with the weak decay of the $b$ quark. by considering the contribution of the resonance $K_{0}^{*}(1430)$, generated from the vector-vector interaction, and the contribution of the intermediate resonance $Y(4710)$, respectively, and try to discuss whether the $Y(4710)$ can be produced in the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ or not according to the LHCb measurements [37], which should be an incentive for the more accurate experimental analysis.
@@ II.1 Contribution of $K_{0}^{*}(1430)$ in the $B^{-}\to J/\psi\eta^{\prime}K^{-}$
Here, we will consider the contribution from the $S$-wave vector meson-vector meson final state interaction and the $B^{-}\to J/\psi VV\to J/\psi\eta^{\prime}K^{-}$ reaction. As shown in Fig. 1, the $b$ quark of the initial $B^{-}$ weakly decays into a $c$ quark and a $W^{-}$ boson, then the $W^{-}$ boson subsequently decays into a $\bar{c}$ quark and an $s$ quark. The $c$ quark and $\bar{c}$ from the $W^{-}$ boson will hadronize into the $J/\psi$ meson. The $s$ quark and the $\bar{u}$ quark from the initial $B^{-}$, together with the quark pair $\bar{q}q=\bar{u}u+\bar{d}d+\bar{s}s$ created from the vacuum with the quantum numbers $J^{PC}=0^{++}$, hadronize into hadron pairs, as follows,
@@
$H$ $=J/\psi\left\{K^{*-}\left(\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}\right)+\bar{K}^{*0}\rho^{-}+K^{*-}\phi\right\}$ (3)
@@
$=$ $-\left[\sqrt{\frac{2}{3}}\left|\bar{K}^{*}\rho,\frac{3}{2},-\frac{1}{2}\right>-\sqrt{\frac{1}{3}}\left|\bar{K}^{*}\rho,\frac{1}{2},-\frac{1}{2}\right>\right]$
@@
This said, diagrammatically we have the mechanisms depicted in Fig. 2, which could prodcue the $K^{*}_{0}(1430)$ through the $S$-wave vector meson-vector meson interaction, and the amplitude could be expressed as,
@@
here we take $\mu=1000$ MeV, and $a=-1.85$ [45, 47]. Since the vector mesons, particularly the $\rho$ and the $K^{*}$, are rather broad, one has to take into account their widths. We convolute the vector-vector $G$ function with the mass distributions of the two vector mesons, i.e., by replacing the $G$ function [58, 54, 55, 57, 56, 59],
[…]
×$\times\int\limits^{(M_{2}+2\Gamma_{2})^{2}}\limits_{(M_{2}-2\Gamma_{2})^{2}}d\tilde{m}^{2}_{2}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{2}-M_{2}^{2}+i\tilde{\Gamma}_{2}\tilde{m}_{2}}$
×$\times G(s,\tilde{m}_{1}^{2},\tilde{m}_{2}^{2})$
@@
[…]
×$\times\int\limits^{(M_{2}+2\Gamma_{2})^{2}}\limits_{(M_{2}-2\Gamma_{2})^{2}}d\tilde{m}^{2}_{2}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{2}-M_{2}^{2}+i\tilde{\Gamma}_{2}\tilde{m}_{2}},$
@@ II.3 Invariant Mass Distribution
where
@@ III Results and Discussions
First, we have fitted to the LHCb invariant mass distributions of $\eta^{\prime}K^{-}$, $J/\psi\eta^{\prime}$, and $J/\psi K^{-}$ without the contribution of $Y(4710)$ [37]. The fitted parameters are $V_{P}=1493.6$, $C=6.9932\times 10^{-5}$, $\phi=-0.13118\pi$, and $\chi^{2}/d.o.f=108.6188/(44-3)=2.6492$. In Fig. 4, the red-dashed curves show the contribution of $\mathcal{T}^{K^{*}_{0}(1430)}$, the blue-dotted curves show the contribution of $\mathcal{T}^{BG}$, and the black solid curves show the contribution of total amplitude. We observe at first glance that the agreement with the three experimental mass distributions is fair. One can see that in the low energy region of the $\eta^{\prime}K^{-}$ mass distribution, this reaction is dominated by the $K_{0}^{*}(1430)$ state, and the interference between $K_{0}^{*}(1430)$ and background plays a key role. Although the basic features of the invariant mass distributions are reproduced, one can find that the peak structure in the high energy region of the $J/\psi\eta^{\prime}$ invariant mass distribution is not well reproduced. Hence, we should consider the a charmonium-like state with mass around 4700 MeV and coupling to $J/\psi\eta^{\prime}$ channel.
@@ IV Conclusions
In this work, we have carried out a detailed study of this process, where we consider the contribution of the $Y(4710)$, and the one of the $K_{0}^{*}(1430)$, dynamically generated from the $S$-wave vector meson-vector meson interaction.
Firstly, we have calculated the invariant mass distributions without the $Y(4710)$, and found that our results cannot describe the peak structure around 4700 MeV in $J/\psi\eta^{\prime}$ invariant mass distribution. By taking account into the contribution of the $Y(4710)$ state, one can find that the peak structure around 4700 MeV in $J/\psi\eta^{\prime}$ mass distribution can be well reproduced. Furthermore, our results of the $\eta^{\prime}K^{-}$ and $J/\psi K^{-}$ invariant mass distributions are in good agreement with the LHCb measurements. Furthermore, we have plotted the Dalitz plot, which is also consistent with the LHCb results.
It it notable that the mass and width of the $Y(4710)$ state have large experimental uncertainties, and this state is not yet listed in RPP. Thus, it is very important to find $Y(4710)$ in more processes. Since the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$ is Okubo-Zweig-Iizuka (OZI) suppressed for the charmonium state, the measurements of the $\mathcal{B}(Y(4710)\to J/\psi\eta^{\prime})$ through the process $B^{-}\to J\psi\eta^{\prime}K^{-}$ or other processes is crucial to understanding the nature of the $Y(4710)$. If the coupling of the $Y(4710)\to J/\psi\eta^{\prime}$ is strong, it implies that the $Y(4710)$ could have the important tetraquark component in addition to the $c\bar{c}$. In all, we advocate that Belle II and LHCb Collaborations could perform the more precise analysis to confirm the evidence of the $Y(4710)$ state [64], which could be helpful to reduce the experimental uncertainties of its mass and width, and to explore its nature.

### v2

@@ Abstract
Motivated by the recent LHCb Collaboration analysis, we have investigated the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ by considering the contributions from the resonances $Y(4710)$ and $K_{0}^{*}(1430)$. Our results are in good agreement with the measured invariant mass distributions and the Dalitz plot of LHCb, which supports the existence of the $K_{0}^{*}(1430)$ and the $Y(4710)$ in the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$. Since the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$ is Okubo-Zweig-Iizuka (OZI) suppressed for the charmonium state, the measurements of the $\mathcal{B}(Y(4710)\to J/\psi\eta^{\prime})$ through the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ or other processes is helpful to understanding the nature of the $Y(4710)$. Thus, we advocate that Belle II and LHCb Collaborations could perform the more precise analysis to confirm the evidence of the $Y(4710)$ state, which could be helpful to reduce the experimental uncertainties of its mass and width, and to explore its nature.
@@ I Introduction
Since the $X(3872)$ has been observed by the Belle Collaboration in 2003 [8], many charmonium-like states have been reported experimentally, and called many attentions [9, 10, 11, 12]. In 2023, the BESIII Collaboration observed a new enhancement structure $Y(4710)$ with a mass of $M=(4704.0\pm 52.3\pm 69.5)$ MeV and a width of $\Gamma=(183.2\pm 114.0\pm 96.1)$ MeV in the cross sections for the process $e^{+}e^{-}\to K_{S}^{0}K_{S}^{0}J/\psi$, and the statistical significance is $4.2\sigma$ [13]. Subsequently, using the data samples with an integrated luminosity of 5.85 $fb^{-1}$ collected at center-of-mass energies from 4.61 to 4.95 GeV, BESIII has measured the cross sections for the process $e^{+}e^{-}\to K^{+}K^{-}J/\psi$, and confirmed the resonance $Y(4710)$ with a significance over $5\sigma$ [14]. Its mass and width are measured to be $M=(4708^{+17}_{-15}\pm 21)$ MeV, $\Gamma=(126^{+27}_{-23}\pm 30)$ MeV, respectively [14], which are consistent with the previous measurements of BESIII within the experimental uncertainties [13].
@@
Although the mass of $Y(4710)$ observed by BESIII is close to the $\psi(5S)$ mass (4704 MeV or 4711 MeV) predicted by the linear potential quark model [15, 16, 17], there are also different theoretical explanations for the structure of the $Y(4710)$ state [18, 19, 20]. In Refs. [21, 22], the authors deem that the $Y(4710)$ is a good candidate of the hidden-charm-hidden-strange tetraquark state $sc]_{\tilde{V}}[\bar{sc}]_{A}-[sc]_{A}[\bar{sc}]_{\tilde{V}}$ with $J^{PC}=1^{--}$ within the framework of the QCD sum rules. In Ref. [23], the authors have studied the charmonium spectrum within an unquenched quark model including coupled channel effects, and found that the $Y(4710)$ can be interpreted as $\psi(4D)$ state. In addition, the $Y(4710)$ can also be interpreted as the hybrid candidates of $H_{1}(4507)$ [24] or $H_{1}[1^{--}](4812)$ [18]. It is notable that the mass and width of the $Y(4710)$ have very large uncertainties, and this state has not yet been listed in Review of Particle Physics (RPP) [25], which implies that the $Y(4710)$ is not well-understood, and still needs confirmation in more processes.
@@
The hadronic decays of the bottom hadrons provide an ideal platform to investigate the charmonium-like states [26, 35, 36, 29, 30, 31, 27, 32, 28, 34, 33]. Recently, the LHCb Collaboration has observed the process $B^{+}\to J/\psi\eta^{\prime}K^{+}$ [37], and given $\mathcal{B}(B^{+}\to J/\psi\eta^{\prime}K^{+})=(3.06\pm 0.29\pm 0.18\pm 0.04)\times 10^{-5}$. From Fig. 2(b) of Ref. [37], one can find a peak structure around 4.7 GeV in the $J/\psi\eta^{\prime}$ invariant mass distribution, close to the mass of $Y(4710)$. If this peak structure can be associated with $Y(4710)$, the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$, which is Okubo-Zweig-Iizuka (OZI) suppressed for the charmonium state decay [38, 39, 40] [^1] In Ref. [40], the selection principle is ‘Among all the possible effective vertices of a given process, the dominant one corresponds to the connected diagram viewed from the composite mode.’ The process $Y(4710)(c\bar{c})\to J/\psi\eta^{\prime}$ is analogous to the disconnected process of Fig. 2(iii) of Ref. [40]., could be helpful to explore its nature, and the precise measurements of this process could be useful to determine the accurate properties of the $Y(4710)$.
@@
In the present work, we will investigate the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ [^2] In this work, we study the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$, the charge conjugation process of the $B^{+}\to J/\psi\eta^{\prime}K^{+}$, because it is convenient to deal with the weak decay of the $b$ quark. by considering the contributions from the resonances $K_{0}^{*}(1430)$ and $Y(4710)$, respectively, and try to discuss whether the $Y(4710)$ can be produced in the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ or not according to the LHCb measurements [37], which should be an incentive for the more accurate experimental analysis.
@@ II.1 Contribution of $K_{0}^{*}(1430)$ in the $B^{-}\to J/\psi\eta^{\prime}K^{-}$
Here, we will consider the contribution from the $S$-wave vector meson-vector meson final state interaction and the $B^{-}\to J/\psi VV\to J/\psi\eta^{\prime}K^{-}$ reaction. As shown in Fig. 1, the $b$ quark of the initial $B^{-}$ weakly decays into a $c$ quark and a $W^{-}$ boson, then the $W^{-}$ boson subsequently decays into a $\bar{c}$ quark and an $s$ quark. The $c$ quark and $\bar{c}$ from the $W^{-}$ boson will hadronize into the $J/\psi$ meson. The $s$ quark and the $\bar{u}$ quark from the initial $B^{-}$, together with the quark pair $\bar{q}q=\bar{u}u+\bar{d}d+\bar{s}s$ created from the vacuum with the quantum numbers $J^{PC}=0^{++}$, hadronize into hadron pairs, as follows, [^3] Here, the produced light quark pairs also can hadronize into other meson pairs, however it is shown that the $K_{0}^{*}(1430)$ could be explained as the dynamically generated state from the vector meson-vector meson interaction in Refs. [45, 46, 47, 48]. Therefore, we have only considered the hadronization of the vector-vector meson pairs here.
@@
$B^{-}$ $=b\bar{u}$ (3)
$\Rightarrow cW^{-}\bar{u}$
$\Rightarrow c\bar{c}s\bar{u}$
$=J/\psi\sum_{i}s\bar{q}_{i}q_{i}\bar{u}$
$=J/\psi\sum_{i}V_{3i}V_{i1}$
$=J/\psi\left\{K^{*-}\left(\frac{\rho^{0}}{\sqrt{2}}+\frac{\omega}{\sqrt{2}}\right)+\bar{K}^{*0}\rho^{-}+K^{*-}\phi\right\}$
@@
[…]
@@
This said, diagrammatically we have the mechanisms depicted in Fig. 2, which could produce the $K^{*}_{0}(1430)$, and the amplitude could be expressed as,
@@
here we take $\mu=1000$ MeV, and a=1.85a=-1.85 [45, 47]. Since the vector mesons, particularly the $\rho$ and the $K^{*}$, are rather broad, one has to take into account their widths. We convolute the vector-vector $G$ function with the mass distributions of the two vector mesons, i.e., by replacing the $G$ function [58, 54, 55, 57, 56, 59],
$\tilde{G}(s)$ $=$ $\frac{1}{N^{2}}\int\limits^{(M_{1}+2\Gamma_{1})^{2}}\limits_{(M_{1}-2\Gamma_{1})^{2}}d\tilde{m}^{2}_{1}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{1}-M_{1}^{2}+i\tilde{\Gamma}_{1}\tilde{m}_{1}}$ (13)
$\times\int\limits^{(M_{2}+2\Gamma_{2})^{2}}\limits_{(M_{2}-2\Gamma_{2})^{2}}d\tilde{m}^{2}_{2}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{2}-M_{2}^{2}+i\tilde{\Gamma}_{2}\tilde{m}_{2}}$
$\times G(s,\tilde{m}_{1}^{2},\tilde{m}_{2}^{2})$
@@
$N^{2}$ $=$ $\int\limits^{(M_{1}+2\Gamma_{1})^{2}}\limits_{(M_{1}-2\Gamma_{1})^{2}}d\tilde{m}^{2}_{1}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{1}-M_{1}^{2}+i\tilde{\Gamma}_{1}\tilde{m}_{1}}$
$\times\int\limits^{(M_{2}+2\Gamma_{2})^{2}}\limits_{(M_{2}-2\Gamma_{2})^{2}}d\tilde{m}^{2}_{2}\left(-\frac{1}{\pi}\right)\mathrm{Im}\frac{1}{\tilde{m}^{2}_{2}-M_{2}^{2}+i\tilde{\Gamma}_{2}\tilde{m}_{2}},$
@@
It is notable that the amplitude of Eq. (8) does not satisfy the unitary requirements. Since we focus the invariant mass distribution and the intermediate resonances in this process, although the unitary condition of the amplitude is not ensured, one will see that our model can give a reasonable description of the experimental data in next section.
@@ II.3 Invariant Mass Distribution
where we introduce the background term from the non-resonant contribution,
@@ III Results and Discussions
First, we have fitted to the LHCb invariant mass distributions of $\eta^{\prime}K^{-}$, $J/\psi\eta^{\prime}$, and $J/\psi K^{-}$ without the contribution of $Y(4710)$ [37]. The fitted parameters are $V_{P}=1493.6$, $C=6.9932\times 10^{-5}$, $\phi=-0.13118\pi$, and $\chi^{2}/d.o.f=108.6188/(44-3)=2.6492$. In Fig. 4, the red-dashed curves show the contribution of $\mathcal{T}^{K^{*}_{0}(1430)}$, the blue-dotted curves show the non-resonant contribution $\mathcal{T}^{BG}$ of Eq. (22), and the black solid curves show the contribution of total amplitude. We observe at first glance that the agreement with the three experimental mass distributions is fair. One can see that in the low energy region of the $\eta^{\prime}K^{-}$ mass distribution, this reaction is dominated by the $K_{0}^{*}(1430)$ state, and the interference between $K_{0}^{*}(1430)$ and background plays a key role. Although the basic features of the invariant mass distributions are reproduced, one can find that the peak structure in the high energy region of the $J/\psi\eta^{\prime}$ invariant mass distribution is not well reproduced. Hence, we should consider the a charmonium-like state with mass around 4700 MeV and coupling to $J/\psi\eta^{\prime}$ channel.
@@ IV Conclusions
In this work, we have carried out a detailed study of this process, where we consider the contributions from the $Y(4710)$ and the $K_{0}^{*}(1430)$. Firstly, we have calculated the invariant mass distributions without the $Y(4710)$, and found that our results cannot describe the peak structure around 4700 MeV in $J/\psi\eta^{\prime}$ invariant mass distribution. By taking account into the contribution of the $Y(4710)$ state, one can find that the peak structure around 4700 MeV in $J/\psi\eta^{\prime}$ mass distribution can be well reproduced. Furthermore, our results of the $\eta^{\prime}K^{-}$ and $J/\psi K^{-}$ invariant mass distributions are in good agreement with the LHCb measurements. Furthermore, we have plotted the Dalitz plot, which is also consistent with the LHCb results.
It it notable that the mass and width of the $Y(4710)$ state have large experimental uncertainties, and this state is not yet listed in RPP. Thus, it is very important to measure the properties of the $Y(4710)$ precisely in more processes. Meanwhile, if the $Y(4710)$ is a charmonium state ($c\bar{c}$), the decay mode of $Y(4710)\to J/\psi\eta^{\prime}$ is OZI-suppressed, and the branching fraction of the mode $\mathcal{B}(Y(4710)\to J/\psi\eta^{\prime})$ is expected to be very small. If the measured branching fraction of this mode through the process $B^{-}\to J/\psi\eta^{\prime}K^{-}$ or other processes is significantly larger than other decay modes, it could imply that the $Y(4710)$ could have the complex components, such as tetraquark component, in addition to the $c\bar{c}$ component. In all, we advocate that Belle II and LHCb Collaborations could perform the more precise analysis to confirm the evidence of the $Y(4710)$ state [64].